Using traditional methods it takes 103 hours to receive an advanced driving license. A new training technique using Computer Aided Instruction (CAI) has been proposed. A researcher used the technique on 50 students and observed that they had a mean of 101 hours. Assume the population variance is known to be 36. Is there evidence at the 0.02 level that the technique performs differently than the traditional method?


The value of the test statistic is −2.36. Find the ?-value of the test statistic. Does it reject the null hypothesis?
Answer #1

Solution:

This is a two tailed test,

P- Value = 2*P(Z < z )

= 2 * P(Z < -2.36 )

= 2 * 0.0091

= 0.0182

The p-value is p = 0.0182, and since p = 0.0182 < 0.02, it is concluded that the null hypothesis is rejected.

There is sufficient evidence at the 0.02 level that the technique performs differently than the traditional method.